# Formal Foundations for Controlled Stochastic Activity Networks

### v1

@@ 5 Stochastic Models
Next, we define a notion of equivalence for Controlled stochastic automata, which is similar to the one proposed for the uncontrolled model [movaghar2001stochastic].
@@
In many realistic applications, the exact computation of optimal policies for continuous-time decision processes are infeasible due to the size and complexity of the underlying state space or stochastic dynamics. This motivates the use of Probably Approximately Correct (PAC) methods [valiant1984theory], which provide theoretical guarantees of $\varepsilon$-optimality with high probability using sampling- or learning-based procedures. Such guarantees make PAC algorithms particularly attractive for systems in which exact analysis is intractable, but near-optimal performance can be learned efficiently.
@@
For any $\varepsilon,\delta>0$, there exist PAC reinforcement learning algorithms that return a policy $\pi$ such that, with probability at least $1-\delta$, the expected performance of $\pi$ is within $\varepsilon$ of the optimal value. Foundational results for finite-state MDPs were established in [kearns2002near, strehl2009reinforcement], with later refinements providing tight bounds for discounted MDPs [lattimore2013pac]. These methods can be adapted to continuous-time decision processes through uniformization or related discretization techniques, extending PAC guarantees to stochastic systems modeled by Controlled SANs.
Integrating PAC guarantees into the framework of Controlled SANs provides a principled way to synthesize near-optimal policies when traditional analytic optimization methods are infeasible. Unlike exact approaches, PAC algorithms offer performance bounds that scales with problem complexity, enabling learning-based control for high-dimensional or partially specified systems [SuttonBarto1998]. This connects the classical theory of optimal and $\varepsilon$-optimal policies [Howard71, puterman1994markov] with modern sample-efficient learning frameworks [valiant1984theory, kearns2002near, strehl2009reinforcement, lattimore2013pac]. This bridge is particularly relevant for AI-driven applications such as autonomous systems, adaptive scheduling, and dependable real-time control, where rigorous performance guarantees and learning-based adaptability must coexist within a unified formal framework.
@@
We examine the synthesis of optimal policies for continuous-time Markov decision processes (CTMDPs) that arise from Controlled SANs when all activities are exponentially timed, assuming finite state and control action sets. By the bounded reward assumptions introduced in the Reward Model and PAC-Optimality Subsection earlier, and standard results for finite-state CTMDPs (see, e.g., [puterman1994markov, feinberg2002handbook]), an optimal stationary deterministic policy is guaranteed to exist. The computational complexity of finding such a policy depends on the specific quantitative objective, as illustrated in the following property:
Proposition 5.18.
@@
Proof Sketch.
The existence of optimal stationary deterministic policies for expected total and discounted rewards in finite-state CTMDPs follows from standard MDP optimality results. For expected total reward, the optimization problem can be formulated as a linear program, guaranteeing a memoryless deterministic optimal policy and polynomial-time solvability. For a discounted reward, the Bellman equations yield a linear system that can be efficiently solved using standard numerical linear algebra methods [puterman1994markov].
The time-bounded reachability (or reward) objective is substantially harder: computing optimal values involves solving systems of ODEs or using uniformization/discretization techniques. The associated decision problems are PSPACE-complete, and optimal strategies may require history-dependent policies [courcoubetis1995complexity]. ∎
@@
Table 3 summarizes the computational complexity and policy characteristics for optimal control in finite-state CTMDPs under the three standard objectives discussed in Proposition 5.18. These results establish that when Controlled SANs yield CTMDPs through exponentially timed activities, existing algorithmic techniques from the CTMDP literature can be directly applied for controller synthesis in safety-critical applications. Such reductions form the basis for extending these methods to more general CSA models, where non-exponential activities lead beyond the classical CTMDP framework.
@@
Expected Total Reward Memoryless deterministic Polynomial time Linear programming [puterman1994markov]
Discounted Reward Memoryless deterministic Polynomial time Linear system or value iteration [puterman1994markov]
Time-Bounded Reachability / Reward May require history dependence PSPACE-complete ODE solving or uniformization [courcoubetis1995complexity]
@@
Proof Sketch.
The proof follows standard arguments for infinite-state MDPs extended to the continuous-time setting by uniformization, combined with bounded reward assumptions and Banach fixed-point arguments. See [puterman1994markov, HernandezLerma1996] for classical existence results for optimal stationary policies. ∎
@@
Having established the existence and computability of optimal policies for both finite- and infinite-state CTMDPs derived from Controlled SANs, we now examine the expressive power of our modeling framework itself. This subsection synthesizes prior results on policy structures, reward semantics, and complexity to show how Controlled Stochastic Automata (CSAs) relate to, subsume, or differ from classical models such as CTMDPs, GSMPs, and other stochastic systems. By clarifying these relationships, we provide a comprehensive perspective on where CSA fits in the hierarchy of stochastic models and what this implies for policy synthesis.
Expressiveness and Optimality of CSA and CTMDP Models
This subsection consolidates the main results on Controlled Stochastic Automata (CSA), connecting the framework introduced earlier in Subsection Policy Types and Subsection Reward Model and PAC-Optimality to the analysis of expressiveness and optimality. We first establish the relationship of CSA to classical models such as CTMDPs and GSMPs, clarifying its expressiveness boundaries. Building on the policy structures and reward accumulation mechanisms defined in Subsections Policy Types and Reward Model and PAC-Optimality, we then present a unified treatment of policy synthesis and value iteration, covering both the general case with arbitrary timing distributions (via GSMP/SMDP embedding) and the special case where exponential timing allows a direct CTMDP reduction.
Proposition 5.22.
Controlled Stochastic Automata (CSA) with general timing distributions strictly generalize classical continuous-time Markov decision processes (CTMDPs) and generalized semi-Markov processes (GSMPs). Specifically:
@@
    Any GSMP can be represented as a CSA by encoding each enabled activity’s clock and transition kernel within the CSA’s timing and control structure.
@@
    CSA subsumes CTMDPs whenever all timing distributions are exponential, since the memoryless property removes the need to track residual holding times.
@@
    CSA cannot, in general, embed stochastic hybrid systems (SHS) or piecewise deterministic Markov processes (PDMPs), because their continuous state dynamics and differential flows cannot be expressed within CSA’s discrete-event semantics.
Sketch of Proof.
For (1), a CSA with general (possibly non-exponential) sojourn times can be mapped to a GSMP by associating each enabled activity $a$ under control $c$ with its timing distribution and transition kernel. The GSMP state, augmented with residual timers, captures exactly the CSA’s event-driven dynamics [Matthes1962, Schassberger1978].
For (2), when all sojourn times are exponential, residual-time information becomes unnecessary due to the memoryless property, and the CSA reduces to a CTMDP. Standard uniformization then yields a discrete-time MDP semantics, enabling classical CTMDP solution techniques.
For (3), SHS and PDMPs include continuous state spaces with deterministic flows governed by differential equations. CSA, being a purely discrete-event formalism with stochastic timing, cannot capture such continuous flows, though it may approximate event-driven subclasses of these models [Cassandras2008, Tomlin1998, Davis1984]. ∎
Proposition 5.23.
Let $V$ be a Controlled Stochastic Automaton (CSA) with bounded reward rates, non-explosive dynamics, and general (not necessarily exponential) timing distributions, as specified under the policy and reward semantics in Subsections Policy Types, and Reward Model and PAC-Optimality. For any discount factor $\beta>0$:
@@
    There exists a sequence of discretized semi-Markov decision processes (SMDPs) approximating $V$, whose optimal value functions converge uniformly to the supremum expected discounted reward achievable under all admissible CSA policies.
@@
    The Bellman operator $T$ associated with each discretized model is a contraction (with modulus $e^{-\beta\Delta}$ for discretization step $\Delta$), so value iteration $V_{k+1}:=TV_{k}$ converges geometrically to the optimal value function $V^{*}_{\Delta}$.
@@
    For any $\varepsilon>0$, a sufficiently fine discretization yields a stationary (possibly randomized) policy that is $\varepsilon$-optimal for $V$.
@@
    When all timing distributions are exponential, the CSA reduces to a CTMDP, which can be uniformized into a discrete-time MDP. In this case, classical CTMDP value iteration applies directly and converges to the optimal value function without state augmentation.
Sketch of Proof. For general timing, we embed $V$ into a GSMP/SMDP by extending each state with residual times of all enabled timed activities, obtaining a Markovian extended state space. Discretizing time with a step size $\Delta>0$, we form the embedded chain, which captures event completions and residual-time evolution. This yields a discrete-time decision process with bounded one-step rewards.
The Bellman operator $T$ for the discretized model is a contraction with modulus $\gamma=e^{-\beta\Delta}$, so the sequence $V_{k+1}:=TV_{k}$ converges geometrically to the unique fixed point $V^{*}_{\Delta}$. Standard error bounds for discretization are well established, both in the SMDP setting (cf. Howard [Howard71, Chapter 15]) and in the MDP setting (cf. Puterman [puterman1994markov, Chapter 7]). These results imply that $V^{*}_{\Delta}$ converges uniformly to the supremum discounted reward $V^{*}$ for the CSA as $\Delta\to 0$. Thus, for any $\varepsilon>0$, a sufficiently fine discretization yields a stationary policy that is $\varepsilon$-optimal for $V$.
Direct uniformization is not applicable unless all sojourn times are exponential. If they are, residual timers are unnecessary by the memoryless property, and $V$ is equivalent to a CTMDP. Standard uniformization transforms this CTMDP into a discrete-time MDP, allowing for the application of classical value iteration, which converges by contraction arguments. ∎
@@
While the above results guarantee the existence of $\varepsilon$-optimal policies and convergence of value iteration, the computational complexity depends heavily on the timing structure. For CTMDPs (exponential case), uniformization leads to tractable dynamic programming with existing solvers. For general CSAs, residual-time augmentation can cause significant state-space growth, and discretization introduces additional overhead. Practical implementations typically rely on truncation, state aggregation, or approximation heuristics to manage this complexity while preserving near-optimality guarantees.
In general, a Controlled Stochastic Automaton (CSA) may include timed activities with arbitrary, non-exponential distributions. Unlike the special case of exponential timing, where the process reduces to a continuous-time Markov decision process (CTMDP) and can be directly uniformized into a discrete-time MDP, non-exponential sojourn times break the memoryless property and preclude such a reduction. To handle the full CSA model, we instead embed it into a generalized semi-Markov process (GSMP) or semi-Markov decision process (SMDP) by augmenting the state with residual timers. This construction preserves the underlying dynamics in a Markovian extended state space and enables discretization-based dynamic programming.
These results collectively show that, while CSA and CTMDP models can exhibit structural complexities that preclude exact optimal policies, $\epsilon$-optimal policies always exist and can be systematically computed. Combined with the Reward Model framework, they ensure that the accumulated reward functional $R_{\tau^{\prime}}^{\tau}(\pi\mid q)$ can be approximated within any tolerance $\epsilon$, allowing robust analysis and synthesis even in complex settings.

### v2

@@ 5 Stochastic Models
Next, we define a notion of equivalence for Controlled stochastic automata, which is similar to the one proposed for the uncontrolled model [49].
@@
In many realistic applications, the exact computation of optimal policies for continuous-time decision processes are infeasible due to the size and complexity of the underlying state space or stochastic dynamics. This motivates the use of Probably Approximately Correct (PAC) methods [71], which provide theoretical guarantees of $\varepsilon$-optimality with high probability using sampling- or learning-based procedures. Such guarantees make PAC algorithms particularly attractive for systems in which exact analysis is intractable, but near-optimal performance can be learned efficiently.
@@
For any $\varepsilon,\delta>0$, there exist PAC reinforcement learning algorithms that return a policy $\pi$ such that, with probability at least $1-\delta$, the expected performance of $\pi$ is within $\varepsilon$ of the optimal value. Foundational results for finite-state MDPs were established in [32, 67], with later refinements providing tight bounds for discounted MDPs [38]. These methods can be adapted to continuous-time decision processes through uniformization or related discretization techniques, extending PAC guarantees to stochastic systems modeled by Controlled SANs.
Integrating PAC guarantees into the framework of Controlled SANs provides a principled way to synthesize near-optimal policies when traditional analytic optimization methods are infeasible. Unlike exact approaches, PAC algorithms offer performance bounds that scales with problem complexity, enabling learning-based control for high-dimensional or partially specified systems [68]. This connects the classical theory of optimal and $\varepsilon$-optimal policies [31, 56] with modern sample-efficient learning frameworks [71, 32, 67, 38]. This bridge is particularly relevant for AI-driven applications such as autonomous systems, adaptive scheduling, and dependable real-time control, where rigorous performance guarantees and learning-based adaptability must coexist within a unified formal framework.
@@
We examine the synthesis of optimal policies for continuous-time Markov decision processes (CTMDPs) that arise from Controlled SANs when all activities are exponentially timed, assuming finite state and control action sets. By the bounded reward assumptions introduced in the Reward Model and PAC-Optimality Subsection earlier, and standard results for finite-state CTMDPs (see, e.g., [56, 20]), an optimal stationary deterministic policy is guaranteed to exist. The computational complexity of finding such a policy depends on the specific quantitative objective, as illustrated in the following property:
Theorem 5.18.
@@
Proof.
We address each part of the theorem separately.
(1) For finite-state CTMDPs with bounded rewards, the expected total reward problem can be formulated as a linear program. This follows from classical results for continuous-time MDPs (see [56]). Linear programming guarantees the existence of an optimal stationary deterministic policy, and the optimization can be performed in time polynomial in the size of the state-action space.
(2) In the discounted setting, the value function satisfies a system of linear equations derived from the Bellman fixed-point operator. Given a discount rate $\beta>0$ and bounded reward functions, the Bellman equations admit a unique solution. The optimal value function can be computed in polynomial time using standard linear algebra techniques (e.g., Gaussian elimination or iterative solvers), and a memoryless deterministic policy achieves the optimal value [56].
(3) This case is substantially more complex. The objective is to maximize the probability (or expected reward) of reaching a target set within a fixed time horizon $T>0$. Unlike the previous two objectives, optimal policies may require history dependence (i.e., memory of past decisions or elapsed time). Numerical methods typically rely on uniformization or discretization, leading to systems of ordinary differential equations (ODEs) whose solution must be approximated. The associated decision problem is known to be PSPACE-complete [51], and the complexity of exact or high-precision computation grows exponentially with the desired numerical accuracy.
@@
Table 3 summarizes the computational complexity and policy characteristics for optimal control in finite-state CTMDPs under the three standard objectives discussed in Theorem 5.18. These results establish that when Controlled SANs yield CTMDPs through exponentially timed activities, existing algorithmic techniques from the CTMDP literature can be directly applied for controller synthesis in safety-critical applications. Such reductions form the basis for extending these methods to more general CSA models, where non-exponential activities lead beyond the classical CTMDP framework.
@@
Expected Total Reward Memoryless deterministic Polynomial time Linear programming [56]
Discounted Reward Memoryless deterministic Polynomial time Linear system or value iteration [56]
Time-Bounded Reachability / Reward May require history dependence PSPACE-complete ODE solving or uniformization [51]
@@
Proof.
We prove each part in turn.
(1) Since both reward components $r^{\prime}$ and $r^{\prime\prime}$ are assumed to be bounded and Borel measurable, the expected total accumulated reward along any feasible policy remains finite as long as the CTMDP is non-explosive. In the case of Controlled SANs, the exponential distributions and structural constraints ensure that the underlying stochastic process is non-explosive. Thus, the value function is well-defined and bounded from above by a finite constant.
(2) We reduce the CTMDP to an equivalent discrete-time MDP using the classical uniformization technique, which transforms the continuous-time process into a discrete-time one with equivalent behavior in terms of expected reward. This technique preserves optimality and allows us to apply well-established results from infinite-state MDP theory. In particular, under bounded reward assumptions and measurability of the transition and reward functions, it is known that an optimal policy exists, possibly requiring randomization or history dependence (see [56, 26]).
(3) When $r^{\prime}$ is continuous and $r^{\prime\prime}$ is weakly continuous (i.e., the expected reward varies continuously with respect to the policy and transition kernel), measurable selection theorems apply. In this case, the supremum in the Bellman functional can be attained by a deterministic stationary policy. Existence follows from measurable selection arguments and upper semi-continuity of the Bellman operator over compact action spaces. Therefore, under these mild continuity assumptions, a stationary deterministic policy attains the optimal value.
When a discount factor $\beta>0$ is introduced, the Bellman operator becomes a contraction in the space of bounded value functions. By the Banach fixed-point theorem, there exists a unique value function, and value iteration converges to it from any initial function. The discounted reward setting thus admits simpler analysis and weaker assumptions, while still ensuring the existence of optimal stationary deterministic policies and convergence of numerical approximation algorithms.
@@
Having established the existence and computability of optimal policies for both finite- and infinite-state CTMDPs derived from Controlled SANs, we now turn to the expressive power of our modeling framework. This subsection synthesizes prior results on policy structures, reward semantics, and complexity to situate Controlled Stochastic Automata (CSA) in the broader landscape of stochastic modeling, including CTMDPs, GSMPs, and GSMDPs. By clarifying these relationships, we aim to offer a comprehensive view of CSA’s expressiveness and implications for policy synthesis.
Expressiveness and Optimality of CSA, CTMDPs, and GSMDPs
This subsection consolidates the main results on Controlled Stochastic Automata (CSA), relating the policy types and reward semantics introduced earlier to the broader spectrum of continuous-time and semi-Markov models. We first formalize the expressiveness hierarchy between CSA and classical models such as CTMDPs, GSMPs, and GSMDPs, and then discuss their relative implications for synthesis and verification.
Theorem 5.22.
Controlled Stochastic Automata (CSA) with general timing distributions strictly generalize continuous-time Markov decision processes (CTMDPs), generalized semi-Markov processes (GSMPs), and even generalized semi-Markov decision processes (GSMDPs). Specifically:
@@
    Any GSMP can be encoded as a CSA by representing concurrently enabled events via local clocks and race-based timing mechanisms.
@@
    CSA subsumes CTMDPs when all timing distributions are exponential, eliminating residual-time tracking via the memoryless property.
@@
    GSMDPs can be simulated by CSAs through explicit representation of general sojourn times and policy-dependent transitions, but CSA provides greater modularity and compositional semantics.
  4. 4.
    CSA cannot, in general, express stochastic hybrid systems (SHS) or piecewise deterministic Markov processes (PDMPs), due to their continuous-state differential dynamics.
Proof.
For (1), GSMPs permit general timing distributions for multiple simultaneously enabled transitions and resolve behavior through the minimum firing time (race semantics). A CSA with general sojourn distributions and gate-based enabling conditions can simulate such behavior by explicitly encoding local clocks and triggering transitions via races. This equivalence is grounded in classical GSMP semantics [42, 61].
For (2), CSA collapses to a CTMDP when all sojourn times are exponential. In this case, the system evolves without residual-time memory, and standard CTMDP uniformization and Bellman equation techniques apply [56].
For (3), GSMDPs extend GSMPs by associating control decisions with transition distributions. CSAs can emulate this behavior by encoding control policies as gate-enabled transitions and representing general distributions via explicit activity semantics. However, CSAs provide additional structure, such as compositional modeling, guard synchronization, and direct alignment with automata-based formal methods, which are absent in classical GSMDP formulations [76].
For (4), stochastic hybrid systems (SHS) and PDMPs allow continuous-valued state variables evolving under differential equations punctuated by stochastic events. CSA models, based on discrete-event formalisms, cannot directly represent continuous-time flows or hybrid transitions [9, 13, 70], though approximations are possible in event-driven subclasses. ∎
Theorem 5.23.
Let $V$ be a Controlled Stochastic Automaton (CSA) with bounded reward rates, non-explosive dynamics, and general (not necessarily exponential) timing distributions, as specified under the policy and reward semantics in Subsections Policy Types and Reward Model and PAC-Optimality. For any discount factor $\beta>0$, the following hold:
@@
    There exists a sequence of discretized generalized semi-Markov decision processes (GSMDPs), equivalently discretized SMDPs, that approximate $V$, whose optimal value functions converge uniformly to the supremum expected discounted reward achievable under all admissible CSA policies.
@@
    The Bellman operator $T$ associated with each discretized model is a contraction with modulus $e^{-\beta\Delta}$ for discretization step $\Delta$, so value iteration $V_{k+1}:=TV_{k}$ converges geometrically to the optimal value function $V^{*}_{\Delta}$.
@@
    For any $\varepsilon>0$, a sufficiently fine discretization yields a stationary (possibly randomized) policy that is $\varepsilon$-optimal for the original CSA $V$.
@@
    When all timing distributions are exponential, the CSA reduces to a CTMDP, which can be uniformized into a discrete-time MDP. In this case, classical CTMDP value iteration applies directly and converges to the optimal value function without residual-time state augmentation.
Proof.
We prove each item in turn.
(1) Consider first the case of arbitrary (non-exponential) timing distributions. Because such distributions violate the memoryless property, the CSA $V$ cannot in general be reduced directly to a CTMDP. Instead, we embed $V$ into a generalized semi-Markov decision process (GSMDP) by augmenting each discrete state with the residual times of all enabled timed activities. Control actions determine which activities are enabled and how their completion events update both the discrete state and the residual timers. This yields a Markovian extended state space faithfully representing the event-driven dynamics of the CSA.
Discretizing time with step size $\Delta>0$ produces a discrete-time SMDP whose optimal value function $V^{*}_{\Delta}$ is well-defined and bounded. Standard results for GSMDPs/SMDPs imply that $V^{*}_{\Delta}\to V^{*}$ uniformly as $\Delta\to 0$, where $V^{*}$ is the supremum expected discounted reward achievable in the CSA.
(2) Let $T$ denote the Bellman operator of the discretized model. Discounting by $\beta>0$ implies that $T$ is a contraction with modulus $\gamma=e^{-\beta\Delta}$. Hence, by the Banach fixed-point theorem, the value iteration sequence $V_{k+1}:=TV_{k}$ converges geometrically to the unique fixed point $V^{*}_{\Delta}$. This fixed point coincides with the optimal value function of the discretized GSMDP (cf. Howard [31, Chapter 15]).
(3) By standard discretization error bounds for semi-Markov decision processes (cf. Puterman [56, Chapter 7]), for any $\varepsilon>0$ there exists a discretization step $\Delta$ sufficiently small such that $\|V^{*}_{\Delta}-V^{*}\|_{\infty}<\varepsilon$. An optimal stationary policy for the discretized model, therefore, induces a stationary (possibly randomized) policy that is $\varepsilon$-optimal for the original CSA $V$.
(4) Finally, when all timing distributions are exponential, the residual-time information is unnecessary due to the memoryless property. In this case, the CSA reduces directly to a CTMDP. Standard uniformization transforms this CTMDP into an equivalent discrete-time MDP, and classical contraction-based value iteration applies directly, converging to the optimal value function without state augmentation. ∎
@@
Although the above results guarantee the existence of $\varepsilon$-optimal policies and convergence of value iteration, the computational complexity depends critically on the timing structure. For exponential timing, uniformization yields tractable dynamic programming with well-established solvers. For general CSAs, the GSMDP embedding requires residual-time augmentation, and discretization may cause substantial state-space growth. Practical implementations, therefore, rely on truncation, abstraction, state aggregation, or approximation heuristics to balance computational tractability with near-optimality guarantees.
In general, a Controlled Stochastic Automaton (CSA) may include timed activities with arbitrary, non-exponential sojourn time distributions. In contrast to the exponential case—where the system dynamics reduce to a continuous-time Markov decision process (CTMDP) and can be uniformly transformed into a discrete-time MDP—the presence of non-exponential timing breaks the memoryless property and thus precludes direct reduction to a CTMDP.
To handle general timing behavior, the CSA can be embedded into a Generalized Semi-Markov Decision Process (GSMDP) by augmenting each state with residual clocks for all concurrently enabled activities. This extended Markovian representation captures the event-driven semantics and enables policy-based analysis. The resulting GSMDP is structurally equivalent to a semi-Markov decision process (SMDP) over an extended state space, where both control actions and general sojourn times are explicitly represented. Such embeddings support dynamic programming via time discretization and residual-timer updates, preserving the accuracy of the underlying stochastic behavior.
This construction enables tractable approximation of optimal control strategies. Specifically, for any discount factor $\beta>0$, there exists a sequence of discretized GSMDPs whose optimal value functions converge uniformly to the supremum expected discounted reward achievable in the original CSA. These results justify the use of discretization-based value iteration, policy improvement, and PAC-style learning even under general timing assumptions.
Combined with our formal reward semantics, these results guarantee that the accumulated reward functional $R_{\tau^{\prime}}^{\tau}(\pi\mid q)$ can be approximated within any desired tolerance $\varepsilon$, enabling robust performance guarantees for decision-making and policy synthesis even in settings with complex timing dynamics.